# Training-free Anomaly Event Detection via LLM-guided Symbolic Pattern Discovery

### v2

@@ Abstract
Anomaly event detection plays a crucial role in various real-world applications. However, current approaches predominantly rely on supervised learning, which faces significant challenges: the requirement for extensive labeled training data and lack of interpretability in decision-making processes. To address these limitations, we present a training-free framework that integrates open-set object detection with symbolic regression, powered by Large Language Models (LLMs) for efficient symbolic pattern discovery. The LLMs guide the symbolic reasoning process, establishing logical relationships between detected entities. Through extensive experiments across multiple domains, our framework demonstrates several key advantages: (i) achieving superior detection accuracy through direct reasoning without any training process; (ii) providing highly interpretable logical expressions that are readily comprehensible to humans; and (iii) requiring minimal annotation effort - approximately 1% of the data needed by traditional training-based methods. To facilitate comprehensive evaluation and future research, we introduce two datasets: a large-scale private dataset containing over 110,000 annotated images covering various anomaly scenarios including construction site safety violations, illegal fishing activities, and industrial hazards, along with a public benchmark dataset of 5,000 samples with detailed anomaly event annotations. Code is available at here.
@@ 1 Introduction
Visual Anomaly Detection (VAD) serves as a cornerstone technology in modern computer vision systems, focusing on identifying irregular patterns or unexpected deviations from established normality in visual data. This field has witnessed exponential growth in its practical applications, playing critical roles in industrial quality control [51, 58], medical diagnosis [54], surveillance systems [32], and autonomous driving safety [5]. The increasing demand for robust anomaly detection solutions has sparked intensive research efforts, leading to significant advancements in bridging the gap between academic innovations and industrial deployments [36, 2, 3, 8, 13, 62, 65, 29]. Despite these advances, developing efficient and reliable VAD systems remains challenging, particularly in scenarios requiring rapid adaptation to new anomaly patterns while maintaining interpretable decision processes.
Figure 1: Illustration of training-based methods alongside out VED-SR pipeline. VED-SR provides low-consumption and plug-and-play solutions.
Despite considerable advancements in Visual Anomaly Detection (VAD), Event-based Anomaly Detection (EAD) involving complex object interactions still presents substantial challenges that have not been fully addressed. There are three main limitations in current methods. First, these methods largely depend on extensive annotated training datasets, which are often limited or extremely costly to acquire in practical settings [19]. This reliance restricts their applicability in various domains. Second, current techniques struggle to grasp the intricate relationships among entities [7]. The significant semantic differences between normal and abnormal behaviors, along with a lack of prior knowledge about anomalies, make it difficult to understand and formulate complex logical rules, particularly in situations with numerous interacting entities or different quantities. Third, the environments in which event detection systems operate, such as processing vast data streams from surveillance cameras and unmanned aerial vehicles [20], present notable scalability issues. The infrequency of anomalies and the marked imbalance between normal and abnormal activities in real-world data further complicate the detection process, rendering traditional training-based methods increasingly ineffective.
Figure 2: The overall performance on the Multi-Event benchmark. Our method outperforms other multimodal fine-tuning methods under multiple abnormal events without requiring training.
To address these challenges, existing approaches can be broadly categorized into two paradigms: implicit modeling and explicit relationship analysis. The implicit modeling approaches aim to capture inter-entity relationships through end-to-end feature learning. GCAD [4] pioneered this direction by employing dual student-teacher architectures for comprehensive anomaly detection. Subsequent works have enhanced this framework: DSKD [67] introduced contextual affinity distillation to better understand normal global context, while GLCF [61] adopted a local-global alignment strategy with a bottleneck architecture. More recently, EfficientAD [1] achieved computational efficiency through knowledge distillation while maintaining relationship modeling capability. The explicit relationship analysis approaches, on the other hand, adopt a more interpretable two-stage strategy: first detecting individual entities, then analyzing their relationships. ComAD [31] exemplifies this direction by combining unsupervised semantic segmentation through K-means clustering on DINO pre-trained features with metrological feature analysis, offering a training-free solution. PSAD [23]further advances this paradigm through few-shot learning, integrating positional features and multiple memory banks for modeling component quantities, arrangements, and fine-grained features. However, while both paradigms have shown promise, they still face fundamental limitations: implicit methods lack interpretability and require substantial training data, while existing explicit methods struggle with complex logical relationships and computational efficiency.
To overcome these limitations, we propose Visual Event Detection with Symbolic Reasoning (VED-SR), a novel training-free framework that synergizes symbolic regression with Large Language Model (LLM) reasoning for anomaly event detection. Our approach fundamentally differs from existing solutions in three key aspects, details in Figure 1. First, rather than relying on direct pattern matching or extensive training, VED-SR leverages LLMs’ symbolic reasoning capabilities and rich commonsense knowledge to construct contextually appropriate anomaly detection models. This enables more nuanced and adaptable anomaly pattern recognition across diverse scenarios. Second, we incorporate symbolic regression, a classical interpretable modeling approach that has demonstrated remarkable effectiveness in capturing intrinsic data patterns while maintaining strong generalization capabilities . The integration of symbolic regression with LLM guidance creates a powerful synergy: the LLM’s prior knowledge steers the symbolic pattern discovery process, while symbolic regression ensures interpretable and generalizable results.
Our framework adopts a training-free, plug-and-play architecture that makes it immediately deployable across different domains without requiring domain-specific training. Extensive experiments across multiple visual event benchmarks, including UCSD Ped2 [52], ERA [35], and our newly introduced Multi-Event and Helmet-Mac datasets, demonstrate the effectiveness of VED-SR. Notably, when integrated with three distinct open-source multimodal object detectors, our method achieves an AUROC of 98.7% on the UCSD Ped2 dataset without any fine-tuning, approaching the state-of-the-art performance of 99.7% achieved by training-based methods. Furthermore, as illustrated in Figure 2, our approach demonstrates comparable and occasionally leading performance on the Multi-Event Dataset. These results, combined with consistent superior performance across other benchmarks, validate the effectiveness and generalizability of our approach. To summarize, Our key contributions include:
  • We propose VED-SR, a novel training-free framework for anomaly event detection that integrates open-set object detection with symbolic regression, guided by Large Language Models. This approach eliminates the need for training data while maintaining competitive detection performance.
  • We develop an efficient symbolic pattern discovery mechanism leveraging LLMs’ reasoning capabilities, which enables the automatic construction of interpretable logical expressions for anomaly detection. This mechanism significantly reduces the annotation effort to approximately 1% of what traditional training-based methods require.
  • We introduce two comprehensive anomaly event detection datasets: a large-scale private dataset containing over 110,000 annotated images covering diverse scenarios (construction site safety violations, illegal fishing activities, and industrial hazards), and a public benchmark dataset of 5,000 samples with detailed annotations to facilitate reproducible research.
  • Extensive experiments demonstrate that our training-free approach achieves comparable or superior performance to state-of-the-art training-based methods across multiple benchmarks (e.g. achieving 98.7% AUROC on the UCSD Ped2 dataset without fine-tuning while demonstrating a remarkable 15% performance improvement on the Helmet-Mac dataset), while providing explicit reasoning paths for human verification.
@@ 2 Related work
2.1 Visual Anomaly Detection
Visual anomaly detection [59] has emerged as a fundamental research direction in computer vision, with various approaches developed to identify deviations from normal patterns in visual data. These approaches can be broadly categorized into several paradigms: memory-based, reconstruction-based, and more recently, vision-language model-based methods. Memory-based approaches, exemplified by PatchCore [41], directly store and utilize normal sample features for anomaly scoring through nearest-neighbor calculations. Reconstruction-based methods, such as DFR [47], employ auto-encoders to rebuild normal features, while knowledge distillation approaches like RD4AD [11] and ViTAD [66] leverage student networks for feature regression. To address the challenge of over-generalization in anomaly scoring, methods like Draem [63] and MRKD [21] introduce synthetic anomalies to enhance model discrimination capabilities. Recent advancements have explored novel directions. RegAD [18] introduces registration-based proxy tasks for representation learning, while GraphCore [58] employs vision isometric invariant graph neural networks for structural feature extraction. The emergence of vision-language models has opened new possibilities, with WinCLIP [19] pioneering the use of CLIP for anomaly detection through text-image similarity computation. This direction has been further enhanced by APRIL-GAN [9] and AnomalyCLIP [68], which address the limitations of manual prompt design and improve model adaptability. Despite these advances [2, 3, 52, 5, 42], the specific domain of anomaly event detection lacks comprehensive datasets and effective approaches for handling complex scenarios. Our work addresses this gap by introducing a dataset focused on safety helmet compliance in construction sites, characterized by complex backgrounds and long-tail distributions of anomalous events
2.2 Prompt Learning
Prompt Learning represents a paradigm shift in utilizing pre-trained models, fundamentally transforming the approach to downstream tasks [53]. Unlike traditional fine-tuning methods, prompt learning aligns downstream tasks with pre-trained models’ inherent capabilities, demonstrating remarkable versatility across various applications. The field’s evolution began with the LAMA dataset [38], which demonstrated the effectiveness of manually crafted cloze templates in probing language model knowledge. Subsequent research has shown that engineered prefix prompts can effectively handle diverse tasks [6], ranging from basic operations to complex reasoning challenges. In the vision-language domain, prompt-based approaches have become increasingly prevalent in anomaly detection applications [43, 45, 44, 48]. Our work extends this paradigm by designing a specialized prompt space for anomaly event detection, leveraging LLMs’ symbolic reasoning capabilities and commonsense knowledge to guide symbolic pattern discovery.
2.3 Open-set Object Detection
The integration of large language models and multimodal learning frameworks has revolutionized open-set object detection [24, 27, 55, 56, 57, 60, 64, 69, 70, 71, 72]. MDETR [22] pioneered this evolution by introducing text-guided visual region alignment through a DETR-like architecture. GenerateU [28] advanced the field by enabling object identification and naming without predefined categories, while GLIP [26] and GroundingDINO [30] demonstrated the effectiveness of combining detection with language grounding. Recent developments include YOLO-World [10], which achieves efficient zero-shot detection through region-text pair pre-training, and APE [46], which unifies detection and grounding capabilities. While these advances have significantly expanded the capabilities of open-set detection, their application to anomaly event detection remained unexplored until our work with VED-SR.
@@ 3 Method
Figure 3: Illustration of the proposed VED-SR. VED-SR mainly consists of logic search and symbolic reasoning. The former module constructs and explores the search space by leveraging structured entity features extracted from an open-set object detector. Th latter module harnesses the symbolic reasoning capabilities of Large Language Models (LLMs) along with their inherent commonsense understanding of anomaly patterns to guide the search process toward more appropriate and rational pathways.
Recent advances in anomaly detection have predominantly focused on developing lightweight and zero-shot detection frameworks. However, achieving training-free operation while maintaining interpretability remains a significant challenge in anomaly event detection. To address this gap, we propose VED-SR, a novel framework that synergistically combines symbolic regression with Large Language Model (LLM) reasoning. Our approach consists of two key components: (1) a symbolic logic search module that discovers interpretable patterns through entity-level analysis, and (2) an automated LLM reasoning module that leverages semantic understanding to guide and accelerate the pattern discovery process. Figure 3 illustrates our framework’s architecture. Formally, given a visual dataset D containing potential anomalous events:
$D=\{x^{n_{1}},x^{n_{2}},...,x^{n_{i}},x^{a_{1}},x^{a_{2}},...,x^{a_{j}}\}$ (1)
where $x^{n_{i}}$ and $x^{a_{j}}$ represent normal and abnormal samples respectively, our goal is to discover a discriminative function $f(\tau)$ that maps input samples to binary anomaly decisions:
$f(\tau):\mathcal{X}\rightarrow\{0,1\}$ (2)
where $\tau$ represents the parameters of the function. Unlike traditional approaches that learn $f(\tau)$ through supervised training, our method discovers interpretable symbolic expressions through an LLM-guided search process. In the following sections, we detail each component of our framework and explain how they work together to achieve efficient and interpretable anomaly detection.
@@ 3.1 Symbolic Logic Search
The core of our anomaly detection framework lies in discovering interpretable symbolic patterns that characterize anomalous events. This section details our symbolic logic search mechanism, which consists of entity extraction, feature representation, and symbolic pattern discovery.
@@
Symbolic Pattern Discovery Given the extracted features, we employ symbolic regression to discover interpretable patterns that distinguish normal from anomalous events. The symbolic regression problem is formulated as:
@@
  1. 1.
    Initializes a population of candidate expressions
  2. 2.
    Evaluates their fitness on the current dataset
  3. 3.
    Applies genetic operators (mutation, crossover) to generate new candidates
  4. 4.
    Selects the best expressions for the next generation
This process generates human-interpretable symbolic expressions that capture meaningful patterns in the data. For example, in a safety helmet detection scenario, a discovered pattern might be:
@@
where $\phi_{p}$, $\phi_{h}$, and $\phi_{d}$ represent the counting functions for persons, helmets, and heads respectively. The discovered patterns serve as initial hypotheses that will be further refined through LLM reasoning, as detailed in the next section.

### v3

# From Objects to Events: Unlocking Complex Visual Understanding in Object Detectors via LLM-guided Symbolic Reasoning

@@ Abstract
Current object detectors excel at entity localization and classification, yet exhibit inherent limitations in event recognition capabilities. This deficiency arises from their architecture’s emphasis on discrete object identification rather than modeling the compositional reasoning, inter-object correlations, and contextual semantics essential for comprehensive event understanding. To address this challenge, we present a novel framework that expands the capability of standard object detectors beyond mere object recognition to complex event understanding through LLM-guided symbolic reasoning. Our key innovation lies in bridging the semantic gap between object detection and event understanding without requiring expensive task-specific training. The proposed plug-and-play framework interfaces with any open-vocabulary detector while extending their inherent capabilities across architectures. At its core, our approach combines (i) a symbolic regression mechanism exploring relationship patterns among detected entities and (ii) a LLM-guided strategically guiding the search toward meaningful expressions. These discovered symbolic rules transform low-level visual perception into interpretable event understanding, providing a transparent reasoning path from objects to events with strong transferability across domains. We compared our training-free framework against specialized event recognition systems across diverse application domains. Experiments demonstrate that our framework enhances multiple object detector architectures to recognize complex events such as illegal fishing activities (75% AUROC, +8.36% improvement), construction safety violations (+15.77%), and abnormal crowd behaviors (+23.16%). Code is available at here.
@@ 1 Introduction
Figure 1: The radar chart at the top illustrates the comparative performance of various models (SymbolicDet(APE [51]), EVA-CLIP [54], DINOv2 [43], LLAVA [33]) across different event detection scenarios. The bottom section provides a visual representation of workplace safety patterns identified through our SymbolicDet framework, showcasing specific conditions like helmet and glove usage in workshop or construction site environments.
Object detection has become a cornerstone of computer vision, enabling machines to identify and locate entities within visual scenes with remarkable accuracy [15, 14, 47, 46, 2, 3, 63, 61, 21, 34, 40]. State-of-the-art detectors can now recognize thousands of object categories with remarkable precision, transforming applications across autonomous driving, surveillance, industrial inspection, and content analysis. However, despite these advances, a fundamental limitation persists: while modern detectors excel at answering “what is present” in an image, they struggle with understanding “what is happening” — the relationships, interactions, and events occurring between detected entities.
Consider a coastal surveillance scenario where a detector identifies “persons” and “fishing rods” with high confidence. Despite perfect detection, the system fails to recognize the critical event of “illegal multi-rod fishing” — where a single individual operates multiple fishing rods, violating conservation regulations. Similarly, in construction site monitoring, a standard detector might accurately identify workers, equipment, and safety gear, yet remain incapable of recognizing the crucial safety violation where a worker operates machinery without proper protective equipment. These limitations stem from the architectural focus of object detectors on identifying discrete entities rather than modeling the compositional logic, relational semantics, and contextual dependencies that define meaningful events. Without the ability to understand “what is happening” beyond “what exists,” even the most accurate detection systems fall short in scenarios requiring nuanced interpretation of object relationships and contextual significance — a fundamental barrier to deploying truly intelligent visual systems in complex real-world scenarios.
Traditionally, bridging this gap has required two unsatisfactory approaches. The first involves developing specialized event recognition systems trained on extensive labeled datasets for each target event, incurring prohibitive annotation costs and limiting generalizability [12, 32, 25]. The second approach employs fine-tuning techniques to adapt object detectors to specific events, which sacrifices their general-purpose nature and still requires substantial task-specific data [35, 49, 16, 64, 26, 55, 62]. These approaches not only demand significant resources but also typically yield black-box models that provide little insight into their reasoning process — a critical limitation in safety-critical or regulated domains where interpretability is essential.
We present a fundamentally different approach that fundamentally reframes the problem: rather than developing specialized event recognition models, we propose to unlock the latent event recognition capabilities inherent in standard object detectors through the integration of LLM-guided symbolic reasoning. Our key insight is that standard object detectors already implicitly encode rich visual information that, when properly interpreted through symbolic reasoning, can reveal complex events and relationships. Rather than treating detectors as mere entity recognizers, we view them as sophisticated visual sensors whose outputs can be transformed into meaningful event understanding through interpretable logical reasoning.
Our framework, consists of three synergistic components. First, an open-vocabulary object detector extracts entity-level information. Second, a symbolic reasoning module discovers logical patterns among these entities through an evolutionary search process, generating human-readable expressions that capture complex relationships. Third, and most innovatively, we leverage Large Language Models (LLMs) to guide this symbolic search, infusing the process with rich world knowledge and semantic understanding that dramatically improves search efficiency and expression quality. This approach offers several significant advantages over existing methods. First, it operates in a training-free manner, requiring no additional labeled data beyond what the underlying detector was trained on. Second, it maintains complete interpretability, with all event recognition decisions expressed as readable logical rules (as exemplified in Figure 1). Finally, our method is detector-agnostic, functioning as a plug-and-play enhancement layer that can augment any object detection system.
Through extensive experiments across multiple datasets, we demonstrate that our approach successfully enhances various detector architectures (APE [51], GLIP [27], and YOLO-World [6]) to recognize complex events including illegal fishing activities, construction safety violations, and abnormal crowd behaviors. In the UCSD Ped2 benchmark [56], our training-free approach achieves 98.7% AUROC, approaching state-of-the-art performance of specialized, training-intensive methods (99.7%), while providing fully transparent reasoning. For safety helmet compliance detection, our method improves recognition accuracy by 15.77% without any domain-specific training. The principal contributions of our work include:
  • We propose a novel framework that unlocks complex event understanding capabilities in standard object detectors through LLM-guided symbolic reasoning, without requiring additional training.
  • We develop an efficient mechanism for discovering interpretable symbolic patterns from object detector, enabling transparent reasoning from object-level detections to event-level understanding.
  • We introduce a structured LLM reasoning process that guides symbolic search, leveraging natural language understanding to discover meaningful patterns while dramatically improving search efficiency.
  • We introduce the Helmet-Mac Dataset, a comprehensive resource containing 12,213 samples specifically designed for construction safety compliance detection, which we make publicly available to the research community.
Through these contributions, we not only enhance the practical utility of object detection systems but also advance our understanding of how compositional reasoning can bridge low-level perception and high-level event semantics. Our work represents a step toward visual AI systems that not only see objects but understand the meaningful events unfolding within visual scenes.
@@ 2 Related work
2.1 Visual Reasoning and Neuro-symbolic
Our work extends beyond conventional object detection to enable reasoning about complex visual events, placing it within the broader visual reasoning paradigm. While we leverage object detector outputs as our foundation, we transform these into symbolic representations suitable for higher-order reasoning. Visual reasoning research has progressed from simple object recognition to complex scene understanding requiring compositional analysis. Traditional approaches have relied on specialized architectures and extensive labeled datasets for specific reasoning tasks. Visual question answering systems interpret images through natural language questions [20, 4, 22, 10, 38, 5], while scene graph generation approaches identify object relationships to construct structured scene representations [23, 28, 57, 31, 60]. However, these methods often lack interpretability or require task-specific training data. Neuro-symbolic models offer a promising direction by combining neural networks’ perceptual strengths with symbolic reasoning’s interpretability and compositionality [11, 1, 39, 52, 59]. These approaches typically extract symbolic representations from visual scenes using neural networks, then apply symbolic reasoning methods to these representations. Our framework advances this paradigm by implementing a novel neuro-symbolic approach where object detectors serve as the neural perception component while a symbolic reasoning layer guided by LLMs performs higher-level event understanding. Unlike traditional implementations requiring custom integration between components, our approach treats existing object detectors as modular perception units, maintaining interpretability while enabling flexible application across visual domains without task-specific training.
2.2 LLMs for Visual Tasks
Our approach uniquely positions LLMs as reasoning guides for symbolic search rather than for direct visual perception. This design allows us to leverage LLMs’ rich world knowledge while maintaining a clear separation between perception (via object detectors) and reasoning (via interpretable symbolic operations). Recent advances in LLMs have demonstrated impressive capabilities in visual understanding [45, 37, 17, 54, 61, 29, 65, 51, 6]. Models such as GPT-4V offer high accuracy in complex visual reasoning with low hallucination rates,making them suitable for complex visual analysis and general-purpose visual AI applications [58]. Despite these advances, directly applying LLMs to visual reasoning presents challenges due to modality gaps and reasoning complexity. Our framework addresses these challenges by using LLMs in their native text domain to guide symbolic pattern discovery over detector outputs. This approach maintains complete interpretability throughout the process — a critical advantage over end-to-end black-box models. By separating perception from reasoning, we combine neural models’ perceptual capabilities with symbolic reasoning’s interpretability, enhanced by LLMs’ semantic understanding, without requiring extensive multimodal training.
2.3 Event Recognition and Understanding
Our work extends into event recognition, where we enable complex visual understanding by reasoning about compositional relationships between detected entities. Traditional event recognition approaches typically rely on specialized architectures trained on event-specific datasets [30, 50, 41, 9]. These methods often struggle with novel event types or complex scenarios requiring compositional reasoning. More recent approaches leveraging large-scale pretraining have improved generalization capabilities but often lack interpretability and explicit reasoning mechanisms. Our framework addresses these limitations by enabling compositional reasoning over object detections to recognize complex events. Our framework addresses these limitations through compositional reasoning over object detections. By transforming detector outputs into symbolic representations, we enable LLM-guided search to identify specific patterns of object interactions characterizing complex events. Unlike methods requiring extensive event-specific training data, our approach can leverage existing object detectors and LLMs’ reasoning capabilities to understand diverse event types without additional visual training.
@@ 3 Method
Figure 2: Illustration of the proposed SymbolicDet. SymbolicDet mainly consists of logic search and symbolic reasoning. The former module constructs and explores the search space by leveraging structured entity features extracted from an open-set object detector. Th latter module harnesses the symbolic reasoning capabilities of Large Language Models (LLMs) along with their inherent commonsense understanding of visual event patterns to guide the search process toward more appropriate and rational pathways.
In practical applications, merely detecting objects often fails to satisfy real-world engineering requirements. Many scenarios demand recognition of complex object relationships or events, which remains a significant challenge in current research. Construction site monitoring requires identifying not just workers and equipment but safety violations; traffic analysis needs to recognize not only vehicles but also dangerous driving patterns; and surveillance systems must detect not merely people but suspicious behaviors. While specialized event recognition systems exist, they typically require extensive training data and lack interpretability. Our framework addresses this challenge by unlocking the latent event understanding capabilities in standard object detectors through LLM-guided symbolic reasoning. Here, we provide a detailed description of our approach, which transforms object detections into interpretable event recognition without additional training. Figure 5 illustrates our framework’s architecture.
Formally, given a visual dataset $\mathcal{D}$ consisting of n images:
$\mathcal{D}=\{\left(I_{1},y_{1}\right),\left(I_{2},y_{2}\right),\ldots,\left(I_{n},y_{n}\right)\}$ (1)
where each pair consists of an image $I_{i}$ and a binary label $y_{i}\in\{0,1\}$ indicating whether a target event $\varepsilon$ occurs in the image. More specifically, for each image $I_{i}$, a standard object detector $D$ produces a set of detections $\mathcal{O}=\{o_{1},o_{2},\ldots,o_{m}\}$, where each detection $o_{j}=(c_{j},b_{j},s_{j})$ consists of a category label $c_{j}$, a bounding box $b_{j}$, and a confidence score $s_{j}$. These detections represent “what exists” in the image. Our approach seeks to discover an interpretable symbolic expression $E$ that operates solely on the detector outputs to recognize the event.
Furthermore, the derived symbolic expression is utilized to assess its capacity in accurately classifying whether the target event is present within an image:
$E:\mathcal{O}_{I}\rightarrow\{0,1\}$ (2)
This symbolic expression effectively bridges the gap between low-level object detections and high-level event understanding, transforming “what exists” into “what is happening” through logical reasoning over detected entities and their relationships.These components work together in a synergistic manner. The overall workflow can be represented as:
$E^{*}=\arg\max_{E\in\mathcal{L}}\mathcal{G}_{LLM}(E,S(E,\mathcal{F}(\mathcal{D})))$ (3)
Where $E^{*}$ is the optimal discovered symbolic expression, $\mathcal{L}$ is the space of all possible expressions in our symbolic language, $\mathcal{F}$ is the object detector, $S$ is a scoring function that evaluates how well an expression distinguishes positive and negative examples, and $\mathcal{G}_{LLM}$ is the LLM guidance mechanism that directs the search toward promising expressions. In the following sections, we detail each component of our framework and how they work together to unlock event understanding capabilities in standard object detectors.
@@ 3.1 Symbolic Logic Search
The core of our framework is the symbolic pattern discovery mechanism that identifies meaningful logical expressions capable of recognizing complex events from object detections. This process begins with extracting structured entity representations from detector outputs and then proceeds to search for effective symbolic patterns.
@@
Symbolic Pattern Discovery Given the entity representation of images, we next seek to discover symbolic patterns that effectively distinguish images containing the target event from those that do not.The symbolic regression problem is formulated as:
@@
where $\phi_{p}$, $\phi_{h}$, and $\phi_{d}$ represent the counting functions for persons, helmets, and heads respectively. While evolutionary search provides a systematic approach to exploring the expression space, its effectiveness is constrained by the stochastic nature of the search process and the exponential growth of the search space with expression complexity. This fundamental challenge highlights the critical role of our LLM guidance mechanism, which strategically directs the evolutionary process toward promising regions of the expression space, balancing exploration with semantic understanding. This synergistic integration, detailed in the following section, enables SymbolicDet to overcome the computational limitations of conventional symbolic approaches while maintaining interpretability.